Samuel has two daughters, Sonya and Alice. One of them is half as old as Sonya. 60 is the total age of both of them. How old is Alice?
Answer by Theo(13342)

s = age of sonya
a = age of alice

total age is 60, therefore s + a = 60

alice is half as old as sonya, therefore a = 1/2 * s

replace a with 1/2 * s in the equation of s + a = 60 to get s + 1/2 * s = 60.

combine like terms to get 3/2 * s = 60

solve for s to get s = 2/3 * 60 = 40

s = 40

a is half of s, therefore 20

40 + 20 = 60

solution looks good.

solution is that sonya is 40 and alice is 20.